A 30.0 g sample of krypton has a temperature of 26 ∘ C at 585 mmHg What is the volume, in milliliters, of the krypton gas?
Answer #1

we know that P = 585mming PV = nRT -0.769 atm V - ART 1 T = (273 +26) = 299K -0.358 X0.0820299 I n = 30 0.769 83.79 2 11.41 L